Trish must be at work exactly 40 hours/week where she earns $10/hour. While at work she can engage in either Hard Work Hours or Easy Work Hours (where Hard Work Hours = 40 – Easy Work Hours). Trish earns a $15/hour commission for each hour of hard work she engages in, and also receives a $200 bonus if she engages in at least 30 hours of hard work in a given week. What is the vertical value of the vertical intercept for Trish’s choice set over (Easy Work, $) bundles?
Answer #1

Maximum total income that Trish can earn = (Wage*Hard work hours) + (Commission*Hard work hours) + Bonus = (10*40) + (15*40) + 200 = 400 + 600 + 200 = $1200

So, the vertical value of the vertical intercept for Trish’s choice set over (Easy Work, $) bundles is $1,200.